OpenStudy (anonymous):

some one please I don't understand this easy problem [(4+3)*5-6]*2 A3 B6 C12 D40

OpenStudy (anonymous):

I got 58

OpenStudy (anonymous):

You are right :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

but there is now answer choice

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

Can you show me the full question?

OpenStudy (anonymous):

that is the full question That is what I don't understand

OpenStudy (anonymous):

Ok do you know what the A3 means?

OpenStudy (anonymous):

It is an answer choice

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ohhh lol ok

OpenStudy (anonymous):

If it is i keep getting 58

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

I really dont know then, maybe we thinking too hard or something. @phi @Preetha @AdamLevine1986

OpenStudy (anonymous):

ok well I gtg If you figure it out just type the answer I will check it later

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

I think you wrote the wrong stuff

OpenStudy (anonymous):

no I wrote every thing right

OpenStudy (texaschic101):

[ (4 + 3 ) * 5 - 6 ] * 2 = [ 7 * 5 - 6 ] * 2 = 35 - 6 * 2 = 29 * 2 = 58